Thirty percent of twice a number is 30. What was the original number?
Found 2 solutions by mananth, Theo:
Answer by mananth(16949)

number be x
30%x= 0.3x
0.3x=30
x=30/0.3
x=100
...
m.ananth@hotmail.ca

Answer by Theo(13342)

x = number

2*x = twice the number

.3*2*x = 30% of it.

.3*2*x = 30

simplify to .6*x = 30

divide both sides by .6 to get x = 30/.6 = 50

confirm by substituting in original equation.

2*50 = 100

100*.3 = 30

answer is confirmed.